OpenStudy (anonymous):

A partial sum of an arithmetic sequence is given. Find the sum. 4 + 9 + 14 + . . . + 484 need help

OpenStudy (anonymous):

The sum forumla you may find useful is \[S_n = \frac{1}{2} \cdot n(a+l) \] You just need to find the number of terms...

OpenStudy (anonymous):

(Where a = first term, l = last term, n = number)

OpenStudy (anonymous):

I cant find n though

OpenStudy (anonymous):

Hint: The nth term = (n-1)d, so 484 = (n-1)*4 => n-1 = 484/4

OpenStudy (anonymous):

sorry, not 4 :/

OpenStudy (anonymous):

a+(n-1)d ***

OpenStudy (anonymous):

=>484 = 4 + (n-1)*5

OpenStudy (anonymous):

so n is 121. because 484/4= 121

OpenStudy (anonymous):

No sorry I misread it, see the post below.

OpenStudy (anonymous):

the post below??

OpenStudy (anonymous):

a+(n-1)d =>484 = 4 + (n-1)*5